one side of a rectangular patio is 4 feet longer than the other.
if the area of the patio is 140 square feet what are the dimentions of the patio?

Answer by user_dude2008(1862)

x(x+4)=140
x^2+4x=140
x^2+4x-140=0
(x-10)(x+14)=0
x=10 or x=-14
Feasible Answer: x=10
Dimensions: 10' by 14'